OpenStudy (anonymous):

Find the value of the variable if P is between J & K JP= 3y + 1 PK= 12y - 4 JK = 75

OpenStudy (anonymous):

you need to draw it out JP=3Y +1 PK=12y-4 JK=75

OpenStudy (anonymous):

on a number line

OpenStudy (anonymous):

if it is two dimensional

OpenStudy (anonymous):

one *

OpenStudy (anonymous):

i am assuming it is on a number line

OpenStudy (anonymous):

JP + PK = JK = 75 3y + 1 + 12y - 4 = 75 15y - 3 = 75 15y = 78 y = 5.2 Plug in y = 5.2, assume that the number line starts at 0 and ends at 75  Since the distance from J to P is going to be the coordinate point of P, P = 16.6 Someone tell me if I did this wrong....